OpenStudy (lgbasallote):

uranium hexafluoride is a solid at room temperature, but it boils at \(56^o C\).determine the density of uranium hexafluoride at \(60^o C\) and 745 torr. i know d = m/v and PV = nRT but how can i use them here? or do i even use them here?

OpenStudy (anonymous):

you can solve for n and get the mass using the periodic table and then you can solve for v

OpenStudy (lgbasallote):

will you please elaborate?

OpenStudy (carniel):

Just get m and v :P

OpenStudy (lgbasallote):

how to get V?

OpenStudy (lgbasallote):

and m

OpenStudy (carniel):

Use the other formula to find them ^_^

OpenStudy (lgbasallote):

how the heck can i get m o.O

OpenStudy (lgbasallote):

and V

OpenStudy (lgbasallote):

i dont think the "other formulas" can be applied

OpenStudy (lgbasallote):

can you demonstrate how>

OpenStudy (nottim):

wats torr

OpenStudy (lgbasallote):

if i remember right 760 torr = 1 atm...torr is a unit of measurement for pressure

OpenStudy (nottim):

p=745 torr v=? n=? R=? T=60-56 (i think) d=? (looking for) m=? v=?

OpenStudy (nottim):

Oh...Waht about Mm?

OpenStudy (nottim):

on what @zbay say.

OpenStudy (nottim):

but you still need m huh

OpenStudy (anonymous):

do you know \[p =\rho \times r \times T \div M\] where rho is density and M is molecular mass

OpenStudy (anonymous):

P is pressure

OpenStudy (anonymous):

since n= m/M , you can get m/V = rho

OpenStudy (lgbasallote):

uhhh could you rewrite that equation you wrote? kinda vague... \[\huge P = \frac{\rho RT}{M}\]

OpenStudy (lgbasallote):

is that right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

looks like it works :)

OpenStudy (lgbasallote):

how did you get this formula again?

OpenStudy (anonymous):

n=m/M right?

OpenStudy (lgbasallote):

oh i see now

OpenStudy (lgbasallote):

so \[\frac{PM}{RT} = \rho\] right? R = 0.0821

OpenStudy (anonymous):

ya exactly

OpenStudy (lgbasallote):

so which T am i going to use?

OpenStudy (lgbasallote):

60?

OpenStudy (nottim):

probably, because that's the one that applies to this situation

OpenStudy (anonymous):

the temperature at which u have to find density

OpenStudy (lgbasallote):

so what's 56 for? just a distraction?

OpenStudy (anonymous):

ya I guess

OpenStudy (lgbasallote):

hmm then i need to convert torr into atm and C into K right

OpenStudy (anonymous):

yup and 1 torr = 1 atm

OpenStudy (lgbasallote):

i thought it was 760 torr = 1 atm o.O

OpenStudy (anonymous):

oo idk

OpenStudy (lgbasallote):

hmm okay thanks

OpenStudy (lgbasallote):

what would be the unit?

OpenStudy (lgbasallote):

i got the answer as 126.76 g/L is that right?

OpenStudy (nottim):

i dunno

OpenStudy (anonymous):

set n to 1 and solve for V then you take the molar mass of the molecule you are looking at and calculate your density

OpenStudy (anonymous):

R is a constant.